OpenStudy (anonymous):

How would you complete the square on the left side of this equation? x^2-6x=15 A. Add 6 to both sides of the equation. B. Subtract 6 from both sides of the equation. C. Add 9 to both sides of the equation. D. Subtract 9 from both sides of the equation.

OpenStudy (welshfella):

you need to convert the left side to (x - 3)^2 if you expand that you'll find that the last number is what you want

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

two steps only what is half of 6?

OpenStudy (anonymous):

3

OpenStudy (misty1212):

right and what is \(3^2\)

OpenStudy (anonymous):

9

OpenStudy (misty1212):

that is your answer , 9

OpenStudy (anonymous):

D

OpenStudy (misty1212):

because if \[x^2-6x=15\] then adding 9 gives \[x^2-6x+9=15+9\\ (x-3)^2=24\]\]

OpenStudy (misty1212):

no dear never subtract when you are completing the square you always ADD

OpenStudy (anonymous):

Thanks (;